I am having a lot of trouble solving this problem. It involves factoring. Here it is:
Factor 1/2x^2+x=1
Thank you for your help.

Answer by AnlytcPhil(1806)

I am having a lot of trouble solving this 
problem. It involves factoring.  Here it is:
Solve by factoring: 

      1/2x² + x = 1

First clear of fractions by multiplying every 
term by LCD = 2

2(1/2x²) + 2(x) = 2(1)

        x² + 2x = 2

    x² + 2x - 2 = 0

That cannot be solved by factoring because the
left side will not factor.  Did you copy the
equation correctly?

Edwin